OpenStudy (anonymous):

Find the average value of f(x)=25−x^2 on the interval [0,1]

hartnn (hartnn):

Average value of the function f(x) over an interval (a,b) is \(\Large \dfrac{1}{(b-a)} \int \limits_a^b f(x)dx\)

hartnn (hartnn):

so, you need to solve, \(\Large \dfrac{1}{(1-0)} \int \limits_0^1 (25-x^2 )dx\)

hartnn (hartnn):

try it out, all you need is \(\int x^n dx = \dfrac{x^{n+1}}{n+1}+c\)

hartnn (hartnn):

why would x be 25 ? x would be x only.

OpenStudy (anonymous):

ok im just confused what to put into the equation

hartnn (hartnn):

have you ever solved an integral before ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

good, so can you solve this, \(\Large \int \limits_0^1 (25-x^2)dx\) \(=\Large \int \limits_0^1 25dx\) \(-\Large \int \limits_0^1 x^2dx\)

OpenStudy (anonymous):

alright 74/3

OpenStudy (anonymous):

Then: Find a value c in the interval [0,1] such that f(c) is equal to the average value

hartnn (hartnn):

so you need f(c) = 74/3 to get f(c), just replace 'x' in f(x) by 'c'

hartnn (hartnn):

f(x) = 25-x^2 f(c) = 25 -c^2 25-c^2 = 74/3 find c from here

OpenStudy (anonymous):

alright thanks

hartnn (hartnn):

welcome ^_^